OpenStudy (voltron21):

evaluate limx->0 (2.5+3.5x)^3 -15.625/ x

ganeshie8 (ganeshie8):

expand the cube

OpenStudy (voltron21):

and then what...

ganeshie8 (ganeshie8):

then u wil see that x cancels out top and bottom

OpenStudy (voltron21):

does it cancel all the x's in the numerator?

ganeshie8 (ganeshie8):

nopes, all we bother about is denominator x here, cuz denominator cant equal 0

ganeshie8 (ganeshie8):

pen down, things wil become clear...

OpenStudy (voltron21):

im still lost

ganeshie8 (ganeshie8):

expand the cube first

OpenStudy (voltron21):

ok i did that but i dont really see where ur going with this

ganeshie8 (ganeshie8):

wat do u get after expanding cube ?

OpenStudy (voltron21):

(2.5+3.5x) three times

OpenStudy (voltron21):

am i supposed to multiply that all together??

ganeshie8 (ganeshie8):

yes, expand it completely use this formula :- \(\large (a+b)^3=a^3+3a^2b+3ab^2+b^3 \)

ganeshie8 (ganeshie8):

\(\large \lim_{x->0} \frac{(2.5+3.5x)^3 -15.625}{x}\) \(\large \lim_{x->0} \frac{(2.5)^3 + 3(2.5)^2(3.5x) + 3(2.5)(3.5x)^2 + (3.5x)^3 -15.625}{x}\) \(\large \lim_{x->0} \frac{15.625 + 3(2.5)^2(3.5x) + 3(2.5)(3.5x)^2 + (3.5x)^3 -15.625}{x}\)

ganeshie8 (ganeshie8):

\(\large \lim_{x->0} \frac{\cancel{15.625} + 3(2.5)^2(3.5x) + 3(2.5)(3.5x)^2 + (3.5x)^3 \cancel{-15.625}}{x}\)

ganeshie8 (ganeshie8):

\(\large \lim_{x->0} \frac{3(2.5)^2(3.5x) + 3(2.5)(3.5x)^2 + (3.5x)^3 }{x}\)

ganeshie8 (ganeshie8):

notice that, you can factor out x from entire numerator, and cancel the denominator

OpenStudy (voltron21):

ok thanks for all the help i got 65.625... wanna help me with 2 more problems?

ganeshie8 (ganeshie8):

\(\large \lim_{x->0} \frac{x [3(2.5)^2(3.5) + 3(2.5)(3.5^2x) + (3.5^3x^2)] }{x}\) \(\large \lim_{x->0} \frac{\cancel{x} [3(2.5)^2(3.5) + 3(2.5)(3.5^2x) + (3.5^3x^2)] }{\cancel{x}}\)

ganeshie8 (ganeshie8):

65.625 is \(\large \color{Red}{\checkmark}\) good work !

ganeshie8 (ganeshie8):

yea sure :)

OpenStudy (voltron21):

youre lovely!

OpenStudy (voltron21):

lim x-0 X^2/sin^2(5x)

ganeshie8 (ganeshie8):

aww thnks :) that keeps me going for next two problems lol

OpenStudy (voltron21):

lol so i changed sin^2 to 1-cos

OpenStudy (voltron21):

but idk what to do now because that doesnt equal 0

ganeshie8 (ganeshie8):

we use this :- \(\large \lim_{x->0} \frac{\sin(x)}{x} = 1\)

ganeshie8 (ganeshie8):

convert given limit to this form

OpenStudy (voltron21):

so the answer would be 1?

ganeshie8 (ganeshie8):

nopes, in the given limit, we dont have same x top and bottom. so you need to rearrange a bit

OpenStudy (voltron21):

so do i have to flip it? like multiply by the inverse?

ganeshie8 (ganeshie8):

Yes

ganeshie8 (ganeshie8):

\(\large \lim_{x->0} \frac{x^2}{\sin^2 5x}\) \(\large \lim_{x->0} (\frac{x}{\sin 5x})^2\) \(\large \lim_{x->0} (\frac{5x}{5\sin 5x})^2\) \(\large \frac{1}{25} \lim_{x->0} (\frac{5x}{\sin 5x})^2\)

ganeshie8 (ganeshie8):

now flip it, and take the limit inside the exponent

ganeshie8 (ganeshie8):

\(\large \frac{1}{25} \lim_{x->0} (\frac{1}{\frac{\sin 5x}{5x}})^2\)

ganeshie8 (ganeshie8):

as x->0, 5x->0

ganeshie8 (ganeshie8):

\(\large \frac{1}{25} (\frac{1}{ \lim_{5x->0} \frac{\sin 5x}{5x}})^2\)

ganeshie8 (ganeshie8):

\(\large \frac{1}{25} (\frac{1}{1})^2\)

OpenStudy (voltron21):

ok so those are essentially go to 0 so the answer is 1/25 got it!

ganeshie8 (ganeshie8):

\(\large \frac{1}{25} \)

OpenStudy (voltron21):

\[f(x)=\sqrt{3x^2+4} / 5x+3\] find hte horizontal asymptote

OpenStudy (voltron21):

\[f(x)= \frac{ \sqrt{3x^2+4} }{ 5x+3 }\]